# Benefits and Pitfalls of Reinforcement Learning for Language Model Planning: A Theoretical Perspective

### v1

@@ Abstract
Recent reinforcement learning (RL) methods have substantially enhanced the planning capabilities of Large Language Models (LLMs), yet the theoretical basis for their effectiveness remains elusive. In this work, we investigate RL’s benefits and limitations through a tractable graph-based abstraction, focusing on policy gradient (PG) and Q-learning methods. Our theoretical analyses reveal that supervised fine-tuning (SFT) may introduce co-occurrence-based spurious solutions, whereas RL achieves correct planning primarily through exploration, underscoring exploration’s role in enabling better generalization. However, we also show that PG suffers from diversity collapse, where output diversity decreases during training and persists even after perfect accuracy is attained. By contrast, Q-learning provides two key advantages: off-policy learning and diversity preservation at convergence. We further demonstrate that careful reward design is necessary to prevent reward hacking in Q-learning. Finally, applying our framework to the real-world planning benchmark Blocksworld, we confirm that these behaviors manifest in practice.
@@ 1 Introduction
We then analyze Q-learning, a paradigm well known in game playing but rarely applied to LLMs (Mnih et al., 2013) (Section 5). Our analysis yields two key findings. First, when trained with only an outcome reward signal, Q-learning suffers from reward hacking; however, incorporating process rewards eliminates this issue. Second, once this issue is addressed, Q-learning offers two theoretical advantages over PG: it converges to a solution that preserves output diversity when achieving optimal training accuracy, and it naturally supports off-policy learning. The latter is particularly important in practice, since rollouts performed with a quantized model or large batch sizes are effectively off-policy, as exemplified by the VeRL framework (Sheng et al., 2024). Finally, we validate all these theoretical findings through experiments.
@@ 2.2 Reinforcement Learning Algorithms
The goal of Q-learning is to approximate the Q-function with the model logits. Let $Q_{\theta}(s_{m},a_{m})$ be the Q-function where $s_{m}=(u_{1},u_{2},\cdots,u_{m})$ is the state, $a_{m}\in\mathcal{V}$ is the action, and $s^{\prime}_{m}=(u_{1},u_{2},\cdots,u_{m},a_{m})$ is their next state. The objective is $\sum_{m}\big(Q_{\theta}(s_{m},a_{m})-[R(s_{m},a_{m})+\max_{a^{\prime}_{m}}Q_{\theta}(s^{\prime}_{m},a^{\prime}_{m})]\big)^{2}$. We denote the logits at step $m$ by $\tilde{\bf u}_{m}$. For an individual trajectory, the loss is given by
@@
$R({\bf u},m)=\begin{cases}\delta_{{\bf u}\in\mathcal{P}}\delta_{u_{m+1}=u_{2}},&\text{If outcome reward},\\[6.0pt] \underbrace{\delta_{u_{m+1}=u_{2}}}_{\text{Target check}}-\underbrace{\delta_{(u_{m},u_{m+1})\not\in\mathcal{E}}}_{\text{Adjacency check}},&\text{If process reward}.\end{cases}$ (3)
@@ 3.2 Characterization of the Stable Point in SFT-Based Learning Dynamics
Building on the observation of Wang et al. (2024b) that next-node prediction depends mainly on the current and target nodes, we adopt the following natural assumption about model expressiveness for our structural characterization. Recall that $u_{2}$ and $u_{m}$ denote the target node and the current node at position $m$, respectively.
@@
The model’s predicted logits for the next token can be expressed as a function of the (target, current) node pair, i.e., there exists a function $\mathbf{f}$ such that the logits $\tilde{\mathbf{u}}_{m}=\mathbf{f}(u_{2},u_{m})$.
Note that in the assumption, $\mathbf{f}$ can be an arbitrary function. We now characterize the structure of the stable point achieved by SFT. Due to space limitations, we defer all the proofs in this paper to the appendix.
@@
Assume Assumption 3.1 holds. Let $N_{u_{2},u_{m},k}$ denote the number of occurrences in the training dataset where the target node is $u_{2}$, the current node is $u_{m}$, and the next node is $k$. The optimal solution of SFT satisfies:
${\exp(\mathbf{f}(u_{2},u_{m})[k])\over\sum_{k^{\prime}}\exp(\mathbf{f}(u_{2},u_{m})[k^{\prime}])}=\frac{N_{u_{2},u_{m},k}}{\sum_{k^{\prime}}N_{u_{2},u_{m},k^{\prime}}}\quad\text{if }\sum_{k^{\prime}}N_{u_{2},u_{m},k^{\prime}}>0.$
If $\sum_{k^{\prime}}N_{u_{2},u_{m},k^{\prime}}=0$, output can be any probability distribution.
@@ 5.1 Theoretical Analysis
At training step $t$, let $i_{t}=u_{2}$, $j_{t}=u_{m}$, $k_{t}$, respectively, denote the target, current, and next nodes. We assume for every $(i,j,k)$, $\exists\underline{N}^{\mathrm{prop}}_{i,j,k}>0$ such that
@@
Assume the RL-training uses the outcome reward $R({\bf u},m)=\delta_{{\bf u}\in\mathcal{P}}\,\delta_{u_{m+1}=u_{2}},$ and a stable point exists under persistent exploration (Assumption 3.1). Then, at any stable point of the Q-learning model, for each fixed target $i$ and $k\neq i$, all logits $\mathbf{f}(i,j)[k]$ take the same value depending only on $i$.
@@
Assume Assumption 3.1 holds and the process reward is used, i.e. $R({\bf u},m)=\delta_{u_{m+1}=u_{2}}-\delta_{(u_{m},u_{m+1})\notin\mathcal{E}}.$ Suppose the score vector $\mathbf{f}(i,j)\in\mathbb{R}^{n}$ is initialized at zero and updated under the persistent exploration assumption with learning rate $\eta$. Then, in the Q-learning model, as $t\to\infty$, $\mathbf{f}^{(t)}(i,j)[i]\longrightarrow{\bf A}[j,i],$ and for $k\neq i$,
@@
Takeaway 5: Different from PG methods, in Q-learning, relying solely on the outcome reward signal can cause reward hacking, whereas introducing process rewards mitigates this issue.
@@
We work under the simplified Transformer setting in Wang et al. (2024b): (1) The token embedding matrix and the output weight matrix are both set to the identity; (2) Attention is fixed entirely on the target node $u_{2}$, so the attention block contributes only the value lookup $\mathbf{W}^{V}[u_{2},\cdot]$; (3) All layer normalizations are removed, and the feedforward block is replaced by a linear map of the form $\mathrm{FFN}({\bf X})={\bf X}{\bf W}^{M}$. Under these assumptions, the logit decomposes as $\tilde{\mathbf{u}}_{m+1}[k]=\mathbf{W}^{M}[u_{m},k]+\mathbf{W}^{V}[u_{2},k],$ where $\mathbf{W}^{M}$ arises from the feed-forward weights and $\mathbf{W}^{V}$ from the value matrix of the attention block.
@@
Theorem 5.1 shows that if only outcome reward is used, the learned logits collapse to a constant across all states for a given target. In contrast, Theorem 5.2 and Theorem 5.3 show that with persistent exploration, process rewards can preserve adjacency and reachability (note that in Theorem 5.3, the constant $c_{k}$ is immaterial in terms of path planning, since for any stable point, $\tilde{\mathbf{u}}_{m+1}[k]=\mathbf{W}^{M}[u_{m},k]+\mathbf{W}^{V}[u_{2},k]=\mathbf{A}[j,k]+\mathbf{R}[i,k]-1$, which is the same as Theorem 5.2). Moreover, the convergence holds even under off-policy sampling, and action diversity is preserved because all feasible next nodes converge to the logit value $1$.
@@ 5.2 Empirical Validations
We first examine the training and test accuracy results in Figure 3(a), where we compare Q-learning under different reward designs and sampling policies. All models are initialized from the same base model. Figure 3(b) states the diversity-accuracy trade-off of Q-learning models, policy gradient models, and the continual SFT model (under different temperatures). Figure 4 illustrates the logits of an on-policy Q-learning model with process rewards and fixed attention (attention fixed on the target node $u_{2}$). The model is initialized from the same base model as in Figure 3, and is further trained with reinforcement steps on all pairs $(s,t)\in\mathcal{D}^{\text{SFT}}$ where $t\in[20]$. In each row $i$, we plot the logits for nodes 02020 (normalized to $[0,1]$) when the current node is 0 and the target node is $i$. White indicates larger logits, black indicates smaller logits, and green frames highlight nodes that are both children of node 0 and ancestors of $i$, corresponding to valid outputs. The empirical results are consistent with the takeaways introduced above, as detailed below.
@@ B.3 Graph Problems with Language Models
There are three major paradigms for analyzing how transformers solve graph-related reasoning tasks. The first is mechanistic interpretability, which reverse-engineers trained model weights (Neel et al., 2023). For example, Cohen et al. (2025) observed that transformers implement a spectral algorithm to compute shortest paths. However, this paradigm largely relies on empirical observation without theoretical grounding. The second paradigm is based on expressiveness analysis (Dai et al., 2024; Sanford et al., 2024; Dai et al., 2025; De Luca & Fountoulakis, 2024), constructing weight configurations that enable transformers to simulate algorithms. Yet, such configurations are often unrealistic for transformers trained via SGD (e.g., embedding vectors explicitly set to $1,2,\dots,L$ (Dai et al., 2024)). The third paradigm investigates gradient dynamics, which is both practical and challenging due to the non-convexity of the optimization landscape. Prior work has analyzed path-finding in directed graphs (Wang et al., 2024b) and compositionality of paths (Zhu et al., 2024).
@@ C.1 Path Planning Algorithm in Transformer
Under the assumption that the output distribution $\hat{\mathbf{u}}_{m}$ depends only on the target node $u_{2}$ and the current node $u_{m}$, we can aggregate identical terms, and express the loss as
$\sum_{u_{2},u_{m}}\left(\sum_{k^{\prime}}N_{u_{2},u_{m},k^{\prime}}\right)\left(\sum_{k}\frac{N_{u_{2},u_{m},k}}{\sum_{k^{\prime}}N_{u_{2},u_{m},k^{\prime}}}\log{\exp(\mathbf{f}(u_{2},u_{m})[k])\over\sum_{k^{\prime}}\exp(\mathbf{f}(u_{2},u_{m})[k^{\prime}])}\right).$
If $\sum_{k^{\prime}}N_{u_{2},u_{m},k^{\prime}}\neq 0$, the expression in brackets is the cross-entropy between the empirical distribution and the distribution of doing softmax on vector $\mathbf{f}(u_{2},u_{m})$. It is minimized when
${\exp(\mathbf{f}(u_{2},u_{m})[k])\over\sum_{k^{\prime}}\exp(\mathbf{f}(u_{2},u_{m})[k^{\prime}])}=\frac{N_{u_{2},u_{m},k}}{\sum_{k^{\prime}}N_{u_{2},u_{m},k^{\prime}}}.$
If $\sum_{k^{\prime}}N_{u_{2},u_{m},k^{\prime}}=0$, the loss does not depend on $\mathbf{f}(u_{2},u_{m})$, so it can be any valid probability distribution. ∎
@@ Appendix D Appendix for Policy Gradient
where $N_{i,j,k}^{R,P,t}$ denote the number of times that $\mathbf{u}[2]=i$, $\mathbf{u}[m]=j$ and $\mathbf{u}[m+1]=k$ for $m\geq 3$ in set $\mathcal{D}^{\text{RL,t}}\cap\mathcal{P}$. Then we can take the gradient and get
@@
where $N_{i,j,k}^{R,t}$ denote the number of times that $\mathbf{u}[2]=i$, $\mathbf{u}[m]=j$ and $\mathbf{u}[m+1]=k$ for $m\geq 3$ in set $\mathcal{D}^{\text{RL,t}}$.

### v2

@@ Abstract
Recent reinforcement learning (RL) methods have substantially enhanced the planning capabilities of Large Language Models (LLMs), yet the theoretical basis for their effectiveness remains elusive. In this work, we investigate RL’s benefits and limitations through a tractable graph-based abstraction, focusing on policy gradient (PG) and Q-learning methods. Our theoretical analyses reveal that supervised fine-tuning (SFT) may introduce co-occurrence-based spurious solutions, whereas RL achieves correct planning primarily through exploration, underscoring exploration’s role in enabling better generalization. However, we also show that PG suffers from diversity collapse, where output diversity decreases during training and persists even after perfect accuracy is attained. By contrast, Q-learning provides two key advantages: off-policy learning and diversity preservation at convergence. We further demonstrate that careful reward design is necessary to prevent Q-value bias in Q-learning. Finally, applying our framework to the real-world planning benchmark Blocksworld, we confirm that these behaviors manifest in practice.
@@ 1 Introduction
We then analyze Q-learning, a paradigm well known in game playing but rarely applied to LLMs (Mnih et al., 2013) (Section 5). Our analysis yields two key findings. First, when trained with only an outcome reward signal, Q-learning suffers from Q-value bias; however, incorporating process rewards eliminates this issue. Second, once this issue is addressed, Q-learning offers two theoretical advantages over PG: it converges to a solution that preserves output diversity when achieving optimal training accuracy, and it naturally supports off-policy learning. The latter is particularly important in practice, since rollouts performed with a quantized model or large batch sizes are effectively off-policy, as exemplified by the VeRL framework (Sheng et al., 2024). Finally, we validate all these theoretical findings through experiments.
@@ 2.1 Path Planning Dataset: Syntax and Data Sources
Running Example (Blocksworld).
To connect this abstraction to real-world LLM planning scenarios, consider the Blocksworld domain (Valmeekam et al., 2023b). In Blocksworld, we are given several colored blocks (e.g., Grey, Black, Red, White) placed either on a table or stacked on each other, and the task is to transform an initial arrangement (source state) into a target arrangement (target state) using valid moves. For example, the source state may place all blocks on the table, and the target state requires stacking them so that Red is on Grey, Grey is on Black, and Black is on White. We map every distinct block configuration to a node in $\mathcal{V}$; edges in $\mathcal{E}$ correspond to valid single moves such as “place White on Grey”. A valid plan is therefore equivalent to a path in $\mathcal{G}$ connecting the source node $s$ and target node $t$.
This abstraction matches natural language planning tasks: in the original benchmark, the LLM is given two textual descriptions of initial and target states and asked to generate a sequence of natural language actions to achieve the goal. In our abstract setup, we strip away language semantics to focus on the core planning structure while retaining the same problem difficulty.
@@ 2.2 Reinforcement Learning Algorithms
The goal of Q-learning is to approximate the Q-function with the model logits. Let $Q_{\theta}(s_{m},a_{m})$ be the Q-function where $s_{m}=(u_{\text{source}},u_{\text{target}},u_{1},\cdots,u_{m})$ is the state, $a_{m}\in\mathcal{V}$ is the action, and $s^{\prime}_{m}=(u_{\text{source}},u_{\text{target}},u_{1},\cdots,u_{m},a_{m})$ is their next state. The objective is $\sum_{m}\big(Q_{\theta}(s_{m},a_{m})-[R(s_{m},a_{m})+\max_{a^{\prime}_{m}}Q_{\theta}(s^{\prime}_{m},a^{\prime}_{m})]\big)^{2}$. We denote the logits at step $m$ by $\tilde{\bf u}_{m}$. For an individual trajectory, the loss is given by
@@
$R({\bf u},m)=\begin{cases}\delta_{{\bf u}\in\mathcal{P}}\delta_{u_{m+1}=u_{\text{target}}},&\text{If outcome reward},\\[6.0pt] \underbrace{\delta_{u_{m+1}=u_{\text{target}}}}_{\text{Target check}}-\underbrace{\delta_{(u_{m},u_{m+1})\not\in\mathcal{E}}}_{\text{Adjacency check}},&\text{If process reward}.\end{cases}$ (3)
@@ 3.2 Characterization of the Stable Point in SFT-Based Learning Dynamics
Building on the observation of Wang et al. (2024b) that next-node prediction depends mainly on the current and target nodes, we adopt the following natural assumption about model expressiveness for our structural characterization. Recall that $u_{\text{target}}$ and $u_{m}$ denote the target node and the current node at position $m$, respectively.
@@
The model’s predicted logits for the next token can be expressed as a function of the (target, current) node pair, i.e., there exists a function $\mathbf{f}$ such that the logits $\tilde{\mathbf{u}}_{m}=\mathbf{f}(u_{\text{target}},u_{m})$.
Note that in the assumption, $\mathbf{f}$ can be an arbitrary function. Our experiments validate this assumption. As shown in Section G.1, the evolution of attention maps during training for SFT, PG, and Q-learning demonstrates that the trained transformer acts primarily as a function of the target and current nodes.
We now characterize the structure of the stable point achieved by SFT. Due to space limitations, we defer all the proofs in this paper to the appendix.
@@
Assume Assumption 3.1 holds. Let $N_{u_{\text{target}},u_{m},k}$ denote the number of occurrences in the training dataset where the target node is $u_{\text{target}}$, the current node is $u_{m}$, and the next node is $k$. The optimal solution of SFT satisfies:
${\exp(\mathbf{f}(u_{\text{target}},u_{m})[k])\over\sum_{k^{\prime}}\exp(\mathbf{f}(u_{\text{target}},u_{m})[k^{\prime}])}=\frac{N_{u_{\text{target}},u_{m},k}}{\sum_{k^{\prime}}N_{u_{\text{target}},u_{m},k^{\prime}}}\quad\text{if }\sum_{k^{\prime}}N_{u_{\text{target}},u_{m},k^{\prime}}>0.$
If $\sum_{k^{\prime}}N_{u_{\text{target}},u_{m},k^{\prime}}=0$, output can be any probability distribution.
@@ 5.1 Theoretical Analysis
At training step $t$, let $i_{t}=u_{\text{target}}$, $j_{t}=u_{m}$, $k_{t}$, respectively, denote the target, current, and next nodes. We assume for every $(i,j,k)$, $\exists\underline{N}^{\mathrm{prop}}_{i,j,k}>0$ such that
@@
Assume the RL-training uses the outcome reward $R({\bf u},m)=\delta_{{\bf u}\in\mathcal{P}}\,\delta_{u_{m+1}=u_{\text{target}}},$ and a stable point exists under persistent exploration (Assumption 5.1) and Assumption 3.1. Then, at any stable point of the Q-learning model, for each fixed target $i$ and $k\neq i$, all logits $\mathbf{f}(i,j)[k]$ take the same value depending only on $i$.
@@
Assume Assumption 3.1 holds and the process reward is used, i.e. $R({\bf u},m)=\delta_{u_{m+1}=u_{\text{target}}}-\delta_{(u_{m},u_{m+1})\notin\mathcal{E}}.$ Suppose the score vector $\mathbf{f}(i,j)\in\mathbb{R}^{n}$ is initialized at zero and updated under the persistent exploration assumption with learning rate $\eta$. Then, in the Q-learning model, as $t\to\infty$, $\mathbf{f}^{(t)}(i,j)[i]\longrightarrow{\bf A}[j,i],$ and for $k\neq i$,
@@
Takeaway 5: Different from PG methods, in Q-learning, relying solely on the outcome reward signal can cause Q-value bias, whereas introducing process rewards mitigates this issue.
@@
We work under the simplified Transformer setting in Wang et al. (2024b): (1) The token embedding matrix and the output weight matrix are both set to the identity; (2) Attention is fixed entirely on the target node $u_{\text{target}}$, so the attention block contributes only the value lookup $\mathbf{W}^{V}[u_{\text{target}},\cdot]$; (3) All layer normalizations are removed, and the feedforward block is replaced by a linear map of the form $\mathrm{FFN}({\bf X})={\bf X}{\bf W}^{M}$. Under these assumptions, the logit decomposes as $\tilde{\mathbf{u}}_{m+1}[k]=\mathbf{W}^{M}[u_{m},k]+\mathbf{W}^{V}[u_{\text{target}},k],$ where $\mathbf{W}^{M}$ arises from the feed-forward weights and $\mathbf{W}^{V}$ from the value matrix of the attention block.
@@
Theorem 5.1 shows that if only outcome reward is used, the learned logits collapse to a constant across all states for a given target. In contrast, Theorem 5.2 and Theorem 5.3 show that with persistent exploration, process rewards can preserve adjacency and reachability (note that in Theorem 5.3, the constant $c_{k}$ is immaterial in terms of path planning, since for any stable point, $\tilde{\mathbf{u}}_{m+1}[k]=\mathbf{W}^{M}[u_{m},k]+\mathbf{W}^{V}[u_{\text{target}},k]=\mathbf{A}[j,k]+\mathbf{R}[i,k]-1$, which is the same as Theorem 5.2). Moreover, the convergence holds even under off-policy sampling, and action diversity is preserved because all feasible next nodes converge to the logit value $1$.
@@ 5.2 Empirical Validations
We first examine the training and test accuracy results in Figure 3(a), where we compare Q-learning under different reward designs and sampling policies. All models are initialized from the same base model. Figure 3(b) states the diversity-accuracy trade-off of Q-learning models, policy gradient models, and the continual SFT model (under different temperatures). Figure 4 illustrates the logits of an on-policy Q-learning model with process rewards and fixed attention (attention fixed on the target node $u_{\text{target}}$). The model is initialized from the same base model as in Figure 3, and is further trained with reinforcement steps on all pairs $(s,t)\in\mathcal{D}^{\text{SFT}}$ where $t\in[20]$. In each row $i$, we plot the logits for nodes 02020 (normalized to $[0,1]$) when the current node is 0 and the target node is $i$. White indicates larger logits, black indicates smaller logits, and green frames highlight nodes that are both children of node 0 and ancestors of $i$, corresponding to valid outputs. The empirical results are consistent with the takeaways introduced above, as detailed below.
@@ B.3 Graph Problems with Language Models
There are three major paradigms for analyzing how transformers solve graph-related reasoning tasks. The first is mechanistic interpretability, which reverse-engineers trained model weights (Neel et al., 2023). For example, Cohen et al. (2025) observed that transformers implement a spectral algorithm to compute shortest paths. However, this paradigm largely relies on empirical observation without theoretical grounding. The second paradigm is based on expressiveness analysis (Dai et al., 2024; Sanford et al., 2024; Dai et al., 2025; De Luca and Fountoulakis, 2024), constructing weight configurations that enable transformers to simulate algorithms. Yet, such configurations are often unrealistic for transformers trained via SGD (e.g., embedding vectors explicitly set to $1,2,\dots,L$ (Dai et al., 2024)). The third paradigm investigates gradient dynamics, which is both practical and challenging due to the non-convexity of the optimization landscape. Prior work has analyzed path-finding in directed graphs (Wang et al., 2024b) and compositionality of paths (Zhu et al., 2024).
@@ C.1 Path Planning Algorithm in Transformer
Under the assumption that the output distribution $\hat{\mathbf{u}}_{m}$ depends only on the target node $u_{\text{target}}$ and the current node $u_{m}$, we can aggregate identical terms, and express the loss as
$\sum_{u_{\text{target}},u_{m}}\left(\sum_{k^{\prime}}N_{u_{\text{target}},u_{m},k^{\prime}}\right)\left(\sum_{k}\frac{N_{u_{\text{target}},u_{m},k}}{\sum_{k^{\prime}}N_{u_{\text{target}},u_{m},k^{\prime}}}\log{\exp(\mathbf{f}(u_{\text{target}},u_{m})[k])\over\sum_{k^{\prime}}\exp(\mathbf{f}(u_{\text{target}},u_{m})[k^{\prime}])}\right).$
If $\sum_{k^{\prime}}N_{u_{\text{target}},u_{m},k^{\prime}}\neq 0$, the expression in brackets is the cross-entropy between the empirical distribution and the distribution of doing softmax on vector $\mathbf{f}(u_{\text{target}},u_{m})$. It is minimized when
${\exp(\mathbf{f}(u_{\text{target}},u_{m})[k])\over\sum_{k^{\prime}}\exp(\mathbf{f}(u_{\text{target}},u_{m})[k^{\prime}])}=\frac{N_{u_{\text{target}},u_{m},k}}{\sum_{k^{\prime}}N_{u_{\text{target}},u_{m},k^{\prime}}}.$
If $\sum_{k^{\prime}}N_{u_{\text{target}},u_{m},k^{\prime}}=0$, the loss does not depend on $\mathbf{f}(u_{\text{target}},u_{m})$, so it can be any valid probability distribution. ∎
@@ Appendix D Appendix for Policy Gradient
where $N_{i,j,k}^{R,P,t}$ denote the number of times that $u_{\text{target}}=i$, $u_{m}=j$ and $u_{m+1}=k$ for $m\geq 1$ in set $\mathcal{D}^{\text{RL,t}}\cap\mathcal{P}$. Then we can take the gradient and get
@@
where $N_{i,j,k}^{R,t}$ denote the number of times that $u_{\text{target}}=i$, $u_{m}=j$ and $u_{m+1}=k$ for $m\geq 1$ in set $\mathcal{D}^{\text{RL,t}}$.